OpenStudy (nthenic_oftime):

Solve the system by the method of your choice. Identify systems with no solution and systems with infinitely many solutions, using set notation to express their solution sets. 4x - 3y = 6 -12x + 9y = -24

OpenStudy (mathstudent55):

Look at both equations. Before you start solving the system, is there a common factor in all terms of the second equation? Is there a common factor in all terms of the second equation?

OpenStudy (nthenic_oftime):

yes @mathstudent55 im sorry i got busy for a moment. first the common factpor of 2 and 3 for the second equation...

OpenStudy (mathstudent55):

The first equation has coefficients 4, -3, and 6. It has no common factors. The second equation has coefficients -12, 9, and -24. -12, 9, and -24 are all divisible by 3. 3 is the common factor for the second equation. We divide every term of the second equation by 3: -4x + 3y = -8 Now we compare it with the first equation. 4x - 3y = 6 Let's multiply the second equation by -1 and write it below the first equation: 4x - 3y = 6 4x - 3y = 8 As you can see, the equations are very similar. The only difference between them is the 6 and the 8.

OpenStudy (mathstudent55):

Now let's try to solve it by the addition method. Multiply the second equation by -1, and write it below the first equation. Then add the equations. 4x - 3y = 6 -4x + 3y = -8 --------------- 0 + 0 = -2 When we add the equations, we get this statement. 0 = -2 Since 0 = -2 is a false statement, that means there is no solution for this system of equations. The solution set is the empty set.